Hi, I have been assigneda rather large piece of homework that is mostly easy, however it does contain one question that is baffling me. I have asked about 5 other people about it before coming here, it is as follows.
John takes a test. He ets 2 points for each right answer and loses 1 point for each one wrong.
Write an equation that he did all 50 problems.
Write an equation showing how he got 40 points
And
Graph out both equations and explain what the point of intersection means.
I have no idea how or where to start on this, and neither do any of my friends or family. Please Help!

Answer by stanbon(75887)

Let "x" be the number of questions where he received 2 points.
Let "y" be the number of questions where he received -1 points.
Then x+y=50
The x answers will gain him 2x points (2 pts. for each x answer)
The y answers will lose him -y points (lose -1 for each y answer)
Then 2x-y=40
You now have two eqations and two unknows:
Add the two equations to get:
3x=90
x=30 (number of questions where he received 2 points)
Substiute that value into the 1st (or the 2nd) equation and solve for "y":
30+y=50
y=20 (number of questions where he lost a point)
Cheers,
Stan H.